Sales were 4,000 at a price of $200 but fell to 3,800 when the price was increased to $220. Calculate the percentage change in sales, the percentage change in price, and the price elasticity of demand.
Answer #1

Here S= 4000 Sa 3800 Pelprice) = 8200 P22 $220 Dt. Change in foire = 220-2007 Xloo - a lot. (increak) Prontage change - firalPED - 4. Change in quentity demord 4. Charge in price - 9,29 & P P2-R = 13800-400g x (2000-20) Yoo 1 220-200 xlo - IPED = 0.5